# reaction term and pure phase initial datum

### v1

@@ Abstract
We consider local and nonlocal Cahn–Hilliard equations with constant mobility and singular potentials including, e.g., the Flory–Huggins potential, subject to no-flux (or periodic) boundary conditions. The main goal is to show that the presence of a suitable class of reaction terms allows to establish the existence of a weak solution to the corresponding initial and boundary value problem even though the initial condition is a pure state. In other words, the separation process takes place even in presence of a pure phase, provided that it is triggered by a convenient reaction term. This fact was already observed by the authors in a previous contribution devoted to a specific biological model. In this context, we generalize the previously-mentioned concept by examining the essential assumptions required for the reaction term to apply the new strategy. Also, we explore the scenario involving the nonlocal Cahn–Hilliard equation and provide illustrative examples that contextualize within our abstract framework.
Keywords: Cahn–Hilliard equation, nonlocal Cahn–Hilliard equation, singular potential, initial pure phase, reaction term, existence of a weak solution, well-posedness.
@@ 1 Introduction
Phase separation is a very relevant phenomenon which appears in a number of different contexts (see, for instance, [19, 21] and references therein). In particular, it has recently become a paradigm in Cell Biology (see, e.g., [6, 18, 43]). Phase separation takes place whenever two (or more) species interact in a way that can be described as a competition between mixing entropy and a demixing effect due to the attraction of particles of the same species. As a prototypical example, one may think of oil and vinegar. From the mathematical viewpoint, a fairly good description of such phenomenon is given by the so-called Cahn–Hilliard equation (see [8, 9], see also [49] and its references). In the case of two species, this equation assumes the following form
@@
where a physically relevant example for the potential density $F$ is given by the classical logarithmic potential (see, e.g., [11] and references therein)
@@
where $0<\vartheta<\vartheta_{0}$. An alternative choice is a nonlocal free energy of the form (see [33, 34], see also [26, 27, 14, 15, 16] and their references)
@@
$(K*v)(x):=\int_{\Omega}K(x-y)v(y)\,{\rm dy,}\quad x\in\Omega.$
@@
In order to ensure mass conservation, the above equations are usually equipped with no-flux boundary conditions on $\Sigma:=\partial\Omega\times(0,T)$, even though periodic boundary conditions can also be considered (see [34]). In the vast mathematical literature devoted to these equations, the basic assumption is always that the spatial average of the initial datum must belong to $(-1,1)$ (see, for instance, [45, Def. 6.1, Thm. 6.1]) that is, the initial datum cannot be a pure phase (i.e., $\varphi=1$ or $\varphi=-1$ on the whole $\Omega$). This is also true in the case of nonconstant and nondegenerate mobility. Instead, in the case of degenerate mobility, pure phases are allowed to be solutions by a suitable splitting of the chemical potential, but no phase separation takes place (see [20], see also [22] and references therein for the nonlocal case). In the context of an isolated system, it is inherent not to start from a pure phase, as such conditions preclude the occurrence of phase separation. Yet, in certain scenarios, phase separation is triggered by an external agent, requiring a pure phase, like $\varphi(0)=-1$ for instance, as initial state (see, e.g., [31, 32]). In such instances, placing constraints on the selection of a pure phase as the initial condition may seem overly restrictive.
Here, inspired by [41], we show that it is indeed possible to establish the existence of a weak solution to Cahn–Hilliard equations (1.4) and (1.5) subject to reaction term and with $F$ like, e.g., (1.2), even though the initial datum is a pure phase. More precisely, letting $d\in\{2,3\}$, we consider the following initial and boundary value problems
@@
where $S(x,\varphi,g)$ is a reaction term and $g$ is a given source, possibly depending on time. Here $\Sigma=\partial\Omega\times(0,T)$. We incidentally notice that in the above system we abuse a bit of notation to denote the function $S(x,\varphi,g)$ by explicitly writing the space variable $x\in\Omega$ which is not specified for all the other variables. As no confusion may arise, that convention will be in force throughout the paper.
The main technical difficulty of the entire work derives from the fact that we can no longer prove that $\mu$ belongs to $L^{2}(0,T;L^{2}(\Omega))$, like in [47, 50] (see also [44, 48] for the nonlocal case with degenerate mobility), through an inequality like [52, Prop. A.2] (see also [35, Sec. 5]). This is due to the fact that here we start with a pure phase. Hence, we have to show first that the spatial average of $\varphi$ instantly detaches from $-1$ (i.e., the phase becomes mixed) and to explicitly compute the behavior close to $-1$ for small times. This is based on ad hoc qualitative analysis of evolution of the spatial average of $\varphi$
@@
The second step consists in generalizing the above mentioned inequality by explicitly tracking the behavior of $F^{\prime}(\varphi)$ close to the pure phase. As we shall see, this argument leads to a local in time $L^{p}$-control in time of $\mu$ for $p\in(1,2)$ which is enough to recover the existence of a weak solution. The fact that we start from $-1$ is actually just a matter of choice. We could also start from $1$ by just taking $-S$ in place of $S$.
The analysis of (1.14) imposes some restrictions on $S$ (see below). Nonetheless, we can handle, for instance, two well-known types of Cahn–Hilliard equations with reaction terms, namely, the so-called Cahn–Hilliard–Oono equation (see [40], see also [37] for theoretical results) and the inpainting model proposed by Bertozzi et al. (see [4, 5], see also [10] for the case of logarithmic potential). The former can be deduced by taking $g\equiv 0$ and setting
$S(x,\varphi,g)=S(\varphi)=-m(\varphi-c)$ (1.15)
where $m>0$ and $c\in(-1,1)$ are given. With reference to the (local) model we recall that $c$ is related to the backward and forward reaction rates of the chemical reaction (cf. [40]). Thus, starting from $-1$ means that the separation process is triggered by the chemical reaction only. Observe that this is an off-critical case since the mass is not conserved. Regarding the latter, we have $g\equiv 0$ and
@@
where $\lambda_{0}>0$, $\chi_{U}$ is the characteristic function of a given measurable set $U$ (the complement of the inpainting region) such that $0<|U|<|\Omega|$ and $c:\Omega\to[-1,1]$ is a given bounded function (i.e., the given image). In this case, taking $-1$ as initial datum means that one can start with a totally white image. Then, the image $\varphi$ evolves keeping itself closer to $c$ in the undamaged region and diffusing into the damaged one.
Originally, reaction terms like (1.15) and (1.16) have been proposed in the local case only. Here we also consider and analyze their nonlocal counterpart. In addition, we present (see Section 5) an application to tumor growth models, where the presence of the source term $g$ enters naturally. We observe that the presence of $g$ in our abstract models is motivated by the fact that very often those system are coupled with other equations which models, for instance, reaction-diffusion processes or account for hydrodynamic effects. Thus, including $g$ is the first step towards possible generalizations of these results to more complicated models.
Here, we prove the existence of a weak solution on a given time interval (see Sections 2 and 3). However, note that, the spatial average of $\varphi$ instantaneously belongs to $(-1,1)$. Hence, all the known results apply. Further issues and possible generalizations to other models will be discussed in Section 6.
@@ 2.1 Notation
$\left\|v-v_{\Omega}\right\|\leq c_{\Omega}\left\|\nabla v\right\|,\quad v\in V,$ (2.1)
@@
$\left\|v\right\|$ $=\mathopen{\langle}{v},{v}\mathclose{\rangle}_{V}^{1/2}=\mathopen{\langle}{{\cal R}v},{{\cal N}v}\mathclose{\rangle}_{V}^{1/2}\leq\left\|v\right\|_{V}^{1/2}\left\|{\cal N}v\right\|_{V}^{1/2}\leq\left\|v\right\|_{V}^{1/2}\left\|v\right\|_{*}^{1/2},\quad v\in V.$ (2.3)
@@ 2.2 Assumptions
    where $m$ is given by (2.17), $c\in L^{\infty}(\Omega)$ is such that $c(x)\in[-1,1]$ for almost every $x\in\Omega$, and $\|c\|_{L^{\infty}(\Omega)}<\frac{|U|}{|\Omega|}$.
@@
Note that the first property (P1) is a classical result in the literature of Cahn–Hilliard problems, and dates back to the ideas in [35, Sec. 5] and [52]: this is typically used to control the $L^{1}$-norm in space of $\beta(\varphi)$ whenever the mass of the initial datum, represented by $r_{0}$ in the above estimates, is confined in the interior of the physical domain $(-1,1)$. Nevertheless, in our case the initial datum is a pure phase so, unfortunately, the first property (P1) cannot be employed as the constants $c_{*}$ and $C_{*}$ strongly depend on $r_{*}$ and $r^{*}$. The main technical novelty that we present is then the refinement of such estimate, i.e., property (P2): here, the qualitative behavior of the constants appearing in (2.18) is explicitly tracked when $r_{0}$ approaches the pure phase $-1$. Note that the constants $c_{\beta},C_{\beta}$ appearing in (2.19) depend only on the threshold value $\delta$, but are independent of $r$ and $r_{0}$. The role of the parameter $\delta$ is only technical as it represents the closeness of $r_{0}$ to the pure phase: for example, one can think to fix $\delta=\frac{1}{3}$. For the corresponding proof, we refer the reader to the Appendix.
We point out that the Flory–Huggins potential (1.2) satisfies A1. Thus, it also enjoys the two properties (P1)(P2). Let us note that Proposition 2.4 also holds in the setting of a three-phases Flory–Huggins potential, for which we refer to [41, Appendix A].
@@ 2.3 Main results
Finally, we provide a continuous dependence estimate provided that $h$ is Lipschitz continuous with respect to the second variable. Namely, we postulate that
$\exists\,L>0:\;|h(x,u_{1},v)-h(x,u_{2},v)|\leq L|u_{1}-u_{2}|,\quad\text{for a.a.~{}}x\in\Omega,$ (2.22)
for all $u_{1},u_{2}\in[-1,1]$ and $v\in{\mathbb{R}}$. Under this additional assumption, we have the following results.
@@
Suppose that A1A3 and (2.22) hold. Let $\varphi_{0j}\in V$ be such that $(\varphi_{0j})_{\Omega}\in[-1,1)$, $j=1,2$ and denote by $\varphi_{j}$ a weak solution originating from $\varphi_{0j}$. Then, there exists $C>0$, depending on $T$, $\Omega$, $F$, $m$, and $h$, such that
@@
Suppose that A1A4 and (2.22) hold. Let $\varphi_{0j}\in H$ be such that $(\varphi_{0j})_{\Omega}\in[-1,1)$, $j=1,2$ and denote by $\varphi_{j}$ a weak solution originating from $\varphi_{0j}$. Then, there exists $C>0$, depending on $T$, $\Omega$, $K$, $F$, $m$, and $h$, such that
@@
Thus, we can conclude that our problems are well posed, provided that we also assume (2.22). This is true, in particular for the Cahn–Hilliard–Oono equation as well as for the inpainting equation (see (1.15) and (1.16), respectively).
@@ 3.1 Step (I)
Thanks to classical results (see, e.g., [50]), the above problem has a weak solution $(\varphi_{\lambda},\mu_{\lambda})$ such that
@@
We can adapt the argument in [23, Sec. 3] for the existence (see also references therein). Summing up, there exists a weak solution $(\varphi_{\lambda},\mu_{\lambda})$ to the above problem, such that
@@ 3.4 Step (IV)
Hence, thanks to the estimate (3.11) we deduce that
@@
Lastly, arguing as in [38, Lemmas 7.3-7.4] and using interpolation, we read (3.2) as a one-parameter family of time-dependent system of elliptic equation with maximal monotone nonlinearity $\beta$ and infer that
@@
Thus, reasoning as above and exploit (3.15), we find that, for any $p\in(1,2)$, there exists $C_{p}>0$ such that
@@ 3.5 Step (V)
Let us stress that this latter holds along a suitable subsequence for every $p\in(1,{\overline{p}})$, in the sense that the choice of the subsequence can be done uniformly with respect to $p\in(1,{\overline{p}})$. This fact can be shown rigorously in the following way. First, by exploiting the estimate (3.32) with the choice $p={\overline{p}}/2$, one can select a suitable subsequence along which it holds that $\beta(\varphi_{\lambda})\to\beta(\varphi)$ weakly in $L^{\frac{{\overline{p}}}{2}}(0,T;H)$. Secondly, let us fix such subsequence and prove that along such fixed subsequence it also holds that $\beta(\varphi_{\lambda})\to\beta(\varphi)$ weakly in $L^{p}(0,T;H)$ for all $p\in(1,{\overline{p}})$. To this end, let $p\in(1,{\overline{p}})$ be arbitrary and let $(\beta(\varphi_{\lambda_{k}}))_{k}$ be an arbitrary subsequence of $(\beta(\varphi_{\lambda}))_{\lambda}$. Thanks to (3.32) it follows immediately that $(\beta(\varphi_{\lambda_{k}}))_{k}$ has a further sub-subsequence $(\beta(\varphi_{\lambda_{k_{j}}}))_{j}$ weakly converging to some limit in $L^{p}(0,T;H)$ as $j\to 0$. As we know already that $\beta(\varphi_{\lambda})\to\beta(\varphi)$ weakly in $L^{\frac{{\overline{p}}}{2}}(0,T;H)$, it necessarily holds that $\beta(\varphi_{\lambda_{k_{j}}})\to\beta(\varphi)$ weakly in $L^{p}(0,T;H)$ as $j\to 0$. This shows that every subsequence of $(\beta(\varphi_{\lambda}))_{\lambda}$ admits a further sub-subsequence weakly converging in $L^{p}(0,T;H)$ to $\beta(\varphi)$: this implies that the whole sequence $(\beta(\varphi_{\lambda}))_{\lambda}$ weakly converges to $\beta(\varphi)$ in $L^{p}(0,T;H)$ as ${\lambda}\to 0$, as desired. As for $\pi({\varphi_{\lambda}})$, we use the above strong convergence of the order parameter, along with its Lipschitz continuity, to infer that

### v2

@@ Abstract
We consider local and nonlocal Cahn–Hilliard equations with constant mobility and singular potentials including, e.g., the Flory–Huggins potential, subject to no-flux (or periodic) boundary conditions. The main goal is to show that the presence of a suitable class of reaction terms allows to establish the existence of a weak solution to the corresponding initial and boundary value problem even though the initial condition is a pure state. This fact was already observed by the authors in a previous contribution devoted to a specific biological model. In this context, we examine the essential assumptions required for the reaction term to ensure the existence of a weak solution. Also, we explore the scenario involving the nonlocal Cahn–Hilliard equation and provide some illustrative examples that contextualize within our abstract framework.
Keywords: Cahn–Hilliard equation, nonlocal Cahn–Hilliard equation, singular potential, initial pure phase, reaction term, existence of a weak solution.
@@ 1 Introduction
Phase separation is a relevant phenomenon which appears in a number of different contexts (see, for instance, [20, 22] and references therein). In particular, it has recently become a paradigm in Cell Biology (see, e.g., [7, 19, 44]). Phase separation takes place whenever two (or more) species interact in a way that can be described as a competition between mixing entropy and a demixing effect due to the attraction of particles of the same species. As a prototypical example, one may think of oil and vinegar. From the mathematical viewpoint, a fairly good description of such phenomenon is given by the so-called Cahn–Hilliard equation (see [8, 9], see also [52] and its references). In the case of two species, this equation assumes the following form
@@
where a physically relevant example of potential density $F$ is given by the logarithmic (a.k.a. the Flory–Huggins) potential (see, e.g., [11] and references therein)
@@
where $0<\vartheta<\vartheta_{0}$. An alternative choice is a nonlocal free energy of the form (see [34, 35], see also [1, 27, 28, 14, 15, 16] and their references)
@@
$(K*v)(x):=\int_{\Omega}K(x-y)v(y)\,{\rm dy,}\quad x\in\Omega,$
@@
To ensure mass conservation, the above equations are usually equipped with no-flux boundary conditions on $\Sigma:=\partial\Omega\times(0,T)$ for the chemical potential $\mu$, even though periodic boundary conditions can also be considered (see, for instance, [35]). In the vast mathematical literature devoted to these equations, in order to prove the existence of a reasonable (say, weak) solution in the case of a singular potential like (1.2), it is always assumed that the spatial average of the initial datum must belong to $(-1,1)$ (see, for instance, [18, Thm. 2.2] and [47, Def. 6.1, Thm. 6.1]). However, it must be noted that, in the case of degenerate mobility, pure phases are allowed to be solutions by a suitable splitting of the chemical potential (see [21], see also [23] and references therein for the nonlocal case). In the case of equation (1.4), this fact is less obvious (see [54, Rem. 4.2]).
Here, inspired by a biological model on the formation of condensates (see [32, 33] and references therein, see also [50]) which has been analyzed in [42], we show that it is indeed possible to establish the existence of a weak solution to Cahn–Hilliard equations (1.4) and (1.5) with $F$ like, e.g., (1.2), equipped with no-flux boundary conditions, and subject to a suitable reaction term, if the initial datum is a pure phase.
More precisely, letting $d\in\{2,3\}$, we consider the following initial and boundary value problems
@@
where $S(x,\varphi,g)$ is a reaction term and $g$ is a given source, possibly depending on time. We incidentally notice that in the above system we abuse a bit of notation to denote the function $S(x,\varphi,g)$ by explicitly writing the space variable $x\in\Omega$ which is not specified for all the other variables. As no confusion may arise, that convention will be in force throughout the paper.
The main technical difficulty of the entire work derives from the fact that we can no longer prove the usual property that $\mu$ belongs to $L^{2}(0,T;L^{2}(\Omega))$, like in [49, 53] (see also [45, 51] for the nonlocal case with degenerate mobility), through an inequality like [54, Prop. A.2] (see also [36, Sec. 5]). This is due to the fact that in our scenario, we start with a pure phase. However, here the mass is not conserved because of the reaction term. Thus, under appropriate assumptions, we can show that the spatial average of $\varphi$ instantly detaches from $-1$ (i.e., the phase becomes mixed) and to explicitly compute the behavior close to $-1$ for small times. This is based on ad hoc qualitative analysis of evolution of the spatial average of $\varphi$ given by
@@
The second step consists in generalizing the above mentioned inequality by explicitly tracking the behavior of $F^{\prime}(\varphi)$ close to the pure phase. As we shall see, this argument leads to a local in time $L^{p}$-control in time of $\mu$ for $p\in(1,2)$ which is enough to recover the existence of a weak solution. The fact that we start from $-1$ is actually just a matter of choice. We could also start from $1$ by just taking $-S$ in place of $S$. This essentially means that we are removing mass instead of adding it.
The analysis of (1.14) imposes some restrictions on $S$ (see below). Nonetheless, we can handle, for instance, two well-known types of Cahn–Hilliard equations with reaction terms, namely, a Cahn–Hilliard–Oono type equation (see [41], see also [38] for theoretical results) and the inpainting model proposed by Bertozzi et al. (see [5, 6], see also [10] for the case of logarithmic potential and [46] for the nonlocal case). The former can be deduced by taking $g\equiv 0$ and setting
$S(x,\varphi,g)=S(\varphi)=-m(\varphi-c),$ (1.15)
where $m>0$ and $c\in(-1,1)$ are given. With reference to the (local) model we recall that $c$ is related to the backward and forward reaction rates of the chemical reaction (cf. [41]). Observe that this is an off-critical case and the mass is not conserved. Regarding the latter, we have $g\equiv 0$ and
@@
where $\lambda_{0}>0$, $\chi_{U}$ is the characteristic function of a given measurable set $U$ (the complement of the inpainting region) such that $0<|U|<|\Omega|$ and $c:\Omega\to[-1,1]$ is a given bounded function (i.e., the given image). In this case, taking $-1$ as initial datum means that one can start with a totally white image. Then, the image $\varphi$ evolves keeping itself closer to $c$ in the undamaged region $U$ and diffusing into the damaged one. However, as we shall see (cf. Rem. 2.2 below), $c$ can only be taken close to ±$\pm 1$ values, provided that $|U|$ is large enough.
Originally, reaction terms like (1.15) and (1.16) have been proposed in the local case only. Here we also consider and analyze their nonlocal counterpart. In addition, in Section 5, we present an application to (solid) tumor growth models, where the presence of the source term $g$ becomes meaningful. This term, in our abstract models, is motivated by the fact that very often the corresponding systems are coupled with other equations (for instance, reaction-diffusion processes). Thus, including $g$ is the first step towards possible generalizations of these results to more refined models.
Here, for a certain class of $S$, we prove the existence of a weak solution on a given time interval and, in the case (1.15), its uniqueness (see Sections 2 and 3). However, note that, the spatial average of $\varphi$ instantaneously belongs to $(-1,1)$. Hence, all the known results apply. Further issues and possible generalizations to other models will be discussed in Section 6.
@@ 2.1 Notation
$\left\|v-v_{\Omega}\right\|\leq c_{\Omega}\left\|\nabla v\right\|,\quad v\in V.$ (2.1)
@@
$\left\|v\right\|$ $=\mathopen{\langle}{v},{v}\mathclose{\rangle}_{V}^{1/2}=\mathopen{\langle}{{\cal R}v},{{\cal N}v}\mathclose{\rangle}_{V}^{1/2}\leq\left\|v\right\|_{V}^{1/2}\left\|{\cal N}v\right\|_{V}^{1/2}\leq\left\|v\right\|_{V}^{1/2}\left\|v\right\|_{*}^{1/2},\quad v\in{V_{0}}.$ (2.3)
@@ 2.2 Assumptions
    where $m$ is given by (2.17), $c\in L^{\infty}(\Omega)$ is such that $c(x)\in[-1,1]$ for almost every $x\in\Omega$, and $\|c\|_{L^{\infty}(\Omega)}<\frac{|U|}{|\Omega|}$. Observe that the latter condition is a bit restrictive since $c$ cannot take ±$\pm 1$ values. However, we can say that the smaller the damaged region is, the closer to ±$\pm 1$ $c$ can be taken. On the other hand, if we consider, for instance, $F$ given by (1.2), then $\varphi$ cannot take the values ±$\pm 1$ either. However, in this case, we can take $\vartheta$ small enough in order to have minima as close as possible to ±$\pm 1$. For further details concerning this issue, see [49, Subsec. 2.1] and the references provided therein.
@@
Note that the first property (P1) is a classical result in the literature of Cahn–Hilliard problems, and dates back to the ideas in [36, Sec. 5] and [54]: this is typically used to control the $L^{1}$-norm in space of $\beta(\varphi)$ whenever the mass of the initial datum, represented by $r_{0}$ in the above estimates, is confined in the interior of the physical domain $(-1,1)$. Nevertheless, in our case the initial datum is a pure phase so, unfortunately, the first property (P1) cannot be employed as the constants $c_{*}$ and $C_{*}$ strongly depend on $r_{*}$ and $r^{*}$. The main technical novelty that we present is then the refinement of such estimate, i.e., property (P2): here, the qualitative behavior of the constants appearing in (2.18) is explicitly tracked when $r_{0}$ approaches the pure phase $-1$. Note that the constants $c_{\beta},C_{\beta}$ appearing in (2.19) depend only on the threshold value $\delta$, but are independent of $r$ and $r_{0}$. The role of the parameter $\delta$ is only technical as it represents the closeness of $r_{0}$ to the pure phase: for example, one can think to fix $\delta=\frac{1}{3}$. For the corresponding proof, we refer the reader to the Appendix.
We point out that the Flory–Huggins potential (1.2) satisfies A1. Thus, it also enjoys the two properties (P1)(P2). Let us note that Proposition 2.4 also holds in the setting of a three-phases Flory–Huggins potential, for which we refer to [42, Appendix A].
@@ 2.3 Main results
Finally, assuming that $h$ is constant, we establish a continuous dependence estimate on the initial datum, which may assume the value $-1$. Extending this result to more general reaction terms remains an open issue. Under this additional assumption, we obtain the following result.
@@
Suppose that A1A3 hold and $h$ is independent of the phase variable $\varphi$. Let $\varphi_{0j}\in V$ be such that $(\varphi_{0j})_{\Omega}\in[-1,1)$, $j=1,2$ and denote by $\varphi_{j}$ a weak solution originating from $\varphi_{0j}$. Then, there exists $C>0$, depending on $T$, $\Omega$, $F$, $m$, and $h$, such that
@@
Suppose that A1A4 hold and $h$ is independent of the phase variable $\varphi$. Let $\varphi_{0j}\in H$ be such that $(\varphi_{0j})_{\Omega}\in[-1,1)$, $j=1,2$ and denote by $\varphi_{j}$ a weak solution originating from $\varphi_{0j}$. Then, there exists $C>0$, depending on $T$, $\Omega$, $K$, $F$, $m$, and $h$, such that
@@
Thus, we can conclude that our problems are well posed, provided that the function $h$ does not depend on the phase variable $\varphi$. This holds, in particular, for the Cahn–Hilliard–Oono equation (see (1.15)), where $h$ is a constant, as well as for the inpainting equation (see (1.16)).
@@ 3.1 Step (I)
Thanks to previous results (see, e.g., [53]), the above problem has a weak solution $(\varphi_{\lambda},\mu_{\lambda})$ such that
@@
We can adapt the argument in [24, Sec. 3] for the existence (see also references therein). Summing up, there exists a weak solution $(\varphi_{\lambda},\mu_{\lambda})$ to the above problem, such that
@@ 3.4 Step (IV)
Here we have used the identity $\int_{\Omega}({\mu_{\lambda}})_{\Omega}\big{(}{\varphi_{\lambda}}-({\varphi_{\lambda}})_{\Omega}\big{)}=({\mu_{\lambda}})_{\Omega}\int_{\Omega}\big{(}{\varphi_{\lambda}}-({\varphi_{\lambda}})_{\Omega}\big{)}=0$ as well as the fact that $\varphi_{0,{\lambda}}$ is constant and, therefore, coincides with its mean value $(\varphi_{0,{\lambda}})_{\Omega}$. Hence, thanks to the estimate (3.11) we deduce that
@@
Lastly, arguing as in [39, Lemmas 7.3-7.4] and using interpolation, we read (3.2) as a one-parameter family of time-dependent system of elliptic equation with maximal monotone nonlinearity $\beta$ and infer that
@@
Thus, reasoning as above and exploiting (3.15), we find that, for any $p\in(1,2)$, there exists $C_{p}>0$ such that
@@ 3.5 Step (V)
Let us stress that this latter holds along a suitable subsequence for every $p\in(1,{\overline{p}})$, in the sense that the choice of the subsequence can be done uniformly with respect to $p\in(1,{\overline{p}})$. This fact can be shown rigorously in the following way. First, by exploiting the estimate (3.32) with $p\in(1,{\overline{p}})$ (e.g., $p=\frac{{\overline{p}}+1}{2}$), one can select a suitable subsequence along which it holds that $\beta(\varphi_{\lambda})\to\beta(\varphi)$ weakly in $L^{\frac{{\overline{p}}}{2}}(0,T;H)$. Secondly, let us fix such subsequence and prove that along such fixed subsequence it also holds that $\beta(\varphi_{\lambda})\to\beta(\varphi)$ weakly in $L^{p}(0,T;H)$ for all $p\in(1,{\overline{p}})$. To this end, let $p\in(1,{\overline{p}})$ be arbitrary and let $(\beta(\varphi_{\lambda_{k}}))_{k}$ be an arbitrary subsequence of $(\beta(\varphi_{\lambda}))_{\lambda}$. Thanks to (3.32) it follows immediately that $(\beta(\varphi_{\lambda_{k}}))_{k}$ has a further sub-subsequence $(\beta(\varphi_{\lambda_{k_{j}}}))_{j}$ weakly converging to some limit in $L^{p}(0,T;H)$ as $j\to 0$. As we know already that $\beta(\varphi_{\lambda})\to\beta(\varphi)$ weakly in $L^{\frac{{\overline{p}}}{2}}(0,T;H)$, it necessarily holds that $\beta(\varphi_{\lambda_{k_{j}}})\to\beta(\varphi)$ weakly in $L^{p}(0,T;H)$ as $j\to 0$. This shows that every subsequence of $(\beta(\varphi_{\lambda}))_{\lambda}$ admits a further sub-subsequence weakly converging in $L^{p}(0,T;H)$ to $\beta(\varphi)$: this implies that the whole sequence $(\beta(\varphi_{\lambda}))_{\lambda}$ weakly converges to $\beta(\varphi)$ in $L^{p}(0,T;H)$ as ${\lambda}\to 0$, as desired. As for $\pi({\varphi_{\lambda}})$, we use the above strong convergence of the order parameter, along with its Lipschitz continuity, to infer that